| Operation Summary for March 12, 2012 - March 14, 2012 |
|
Monday March 12 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest, and FTBF T-1008 experiments taking beam. During the day shift, Operations turned off beam to SeaQuest and to the Neutrino Area Switchyard (SY) enclosures, N01/NM1, NM3, and NM4 for an access at 11:06 AM. Controls personnel worked on a SY CAMAC crate during this downtime. MI personnel began a g-2 (g minus 2) study at 1:11 PM. Operations resumed sending beam to SeaQuest around 1:25 PM. During the evening shift, the MI g-2 study ended at 5:53 PM. At 10:05 PM, a Linac RF station (LRF1) tripped off for the first of three times in about 90 minutes; it reset. Tuesday March 13 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1008. During the midnight shift, Operations prepared the Neutrino Area enclosures for access at 7:28 AM. At 7:50 AM, Neutron Therapy Facility (NTF) personnel prepared to treat one patent later in the morning. During the day shift, FESS personnel reported at 8:17 AM that they would be working on the CUB 3000 gallon fill system. Operations prepared SeaQuest for access at 8:27 AM. Operations reported at 12:01 PM that Recycler beam operations were completed; the beam was dumped and the beam permits were pulled. This marked the end of Proton beam in the Recycler. A Booster person began a study at 1:07 PM. The G2 access ended and operations established beam to SeaQuest at 1:57 PM. During the evening shift, at 4:47 PM, the NuMI beam permit tripped off for the first of five times in four hours due to a power supply (E:V118) tripping off; everything reset. At 5:17 PM, a SY power supply (S:V424) tripped off and would not reset; Operators investigated, smelled burnt electronics, and called in a power supply person who repaired a failed contactor. Operators added LCW to the water skid for a LRF5 at 7:43 PM. Wednesday March 14 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1008. During the midnight shift, at 3:25 AM, the NuMI beam permit tripped off for the first of two times in about three hours due to E:V118 tripping off; everything reset. Operations turned off beam to SeaQuest at 7:31 AM and Operators prepared G2, N01, NM2, NM3, and NM4 for access. Operations Coordinator (OpsCo) Plans for Today The OpsCo agrees with the Booster and Muon Ring study plans for today. He also agrees with Linac’s plan to work on LRF1, with MI plans to switch to higher capacity pond water pumps, and with a Controls plan to turn off a Tevatron Low Level RF crate. He will look for the right time to allow Water Group personnel to search for the LCW leak in the MI-62 LCW system. SeaQuest hopes to come back on at 4 PM. FTBF experiment T-1008 completed its run today. |
|
More Information Comments and Suggestions |
| last modified 03/14/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |